OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (−1, 6) and has a slope of −3. y − 6 = −3x − 3 y − 6 = −3(x + 1) y = −3x + 3 3x + y = 3

OpenStudy (anonymous):

The equation we'll be using here is: \[y-y_1=m(x-x_1)\]Where m is the slope of the line, and x_1,y_1 are the coordinates of any point on the line. We have the slope and a point on the line, so you can plug in and solve.

OpenStudy (anonymous):

@vinnv226 so it would be B? y − 6 = −3(x + 1)?

OpenStudy (anonymous):

Yes, exactly right

OpenStudy (anonymous):

@vinnv226 THANKS! :)

OpenStudy (anonymous):

No problem, glad I could help